# and Qian Liu

### v1

@@ III-E.2 Complexity Analysis
Next, we provide a brief analysis of the algorithmic complexity. The primary computational complexity lies in calculating the EM domain pattern selection matrix $\mathbf{S}_{\mathrm{EM}}$ in (36) over the long-timescale, and determining the phase-shift selection matrix $\mathbf{S}_{\mathrm{RF}}$ in (52) over the medium-timescale. Specifically, the computational complexity of solving problem (36) in each super-frame over the long-timescale, which involves an $N_{\mathrm{t}}P$-dimensional variable and requires handling $2N_{\mathrm{t}}P$ second-order cone (SOC) constraints, is of order $\mathcal{O}\{\sqrt{2N_{\mathrm{t}}P+N_{\mathrm{t}}+1}N_{\mathrm{t}}P(N_{\mathrm{t}}^{2}P^{2}+2N_{\mathrm{t}}P+N_{\mathrm{t}})\}$ [34]. Similarly, the computational complexity of solving problem (52) in each frame during the medium-timescale, which entails calculating an $N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}$-dimensional variable along with $2N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}$ SOC constraints, is of order $\mathcal{O}\{\sqrt{2N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}+N_{\mathrm{t}}+1}N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}(N_{\mathrm{t}}^{2}N_{\mathrm{RF}}^{2}2^{2B}+2N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}+N_{\mathrm{t}})\}$.
@@ III-E.3 Pilot Overhead Analysis
The pilot overhead in the proposed scheme primarily stems from channel sampling at the medium-timescale and short-timescale stages. In the medium-timescale, CS-based channel estimation requires $N_{\mathrm{Beam}}K/N_{\mathrm{RF}}$ pilot symbols to obtain the RF domain high-dimensional statistical channel sample $\mathbf{h}_{\mathrm{RF},k}^{(t_{\mathrm{M}})},\forall k$, as described in Section III-C.In the short-timescale, $N_{\mathrm{RF}}K$ pilot symbols are needed for estimating the BB domain low-dimensional effective channel $\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})}$, as described in Section III-D. Therefore, the total pilot overhead across the channel statistics coherence time is expressed as $(N_{\mathrm{Beam}}K/N_{\mathrm{RF}}+N_{\mathrm{RF}}KT_{\mathrm{S}})T_{\mathrm{M}}T_{\mathrm{L}}$. This is notably less than that of the real-timescale beamforming approaches, which requires $(N_{\mathrm{Beam}}K/N_{\mathrm{RF}}+N_{\mathrm{RF}}K)T_{\mathrm{S}}T_{\mathrm{M}}T_{\mathrm{L}}$.
@@ IV Simulation Results
In this section, we provide simulation results to demonstrate the effectiveness of the proposed tri-timescale tri-hybrid beamforming design algorithm with reconfigurable antennas in enhancing communication performance while ensuring hardware efficiency. Unless otherwise specified, the BS is equipped with $N_{\mathrm{t}}=32$ antennas and the center carrier frequency is set to $f_{\mathrm{c}}=28$ GHz. The low-resolution PS used in the RF domain has quantized phase controlled by $B=3$ bits. Without loss of generality, we assume that the number of UEs $K$ is equal to the number of RF chains $N_{\mathrm{RF}}$. In the EM domain, each pixel-based reconfigurable antenna enables the generation of $P=7$ distinct directional radiation patterns by adjusting the electronic switches between pixels [8], with $M=180$ uniform sampling points in the azimuth dimension. In this scenario, there are $K=2$ UEs, all operating under a scattering cluster channel with $C=2$ clusters, each of which consists of $L_{c}=3$ multi-path rays. The nominal AoDs $\overline{\theta}_{k,c}$ of these scattering clusters are randomly distributed within $[-\pi/3,\pi/3]$, while the angle spread is set to $\varsigma_{k,c}=\pi/36$. The path loss is expressed as $\alpha_{k,c,l_{c}}=10^{-\frac{C_{0}}{10}}(\frac{r_{k,c,l_{c}}}{D_{0}})^{-\kappa_{k,c,l_{c}}}$ with $\kappa_{k,c,l_{c}}\in[2.5,3]$ being the attenuation factor, $C_{0}=30$ dB, $D_{0}=1$m. The propagation distance $r_{k,c,l_{c}}$ is composed of the nominal distance $\overline{r}_{k,c}\in[50\mathrm{m},100\mathrm{m}]$ and $\iota=0.5\mathrm{m}$. The noise power of UEs is set as $\sigma_{k}^{2}=-70$dBm. The tri-timescale framework is structured with $T_{\mathrm{L}}=10$ super-frames for the long-timescale, $T_{\mathrm{M}}=50$ frames for the medium-timescale, and $T_{\mathrm{S}}=200$ time-slots for the short-timescale.
@@
Figure 5: Convergence performance of the long-timescale radiation beamformer design and medium-timescale analog beamformer design.
@@
Figure 6: The sum-rate versus the transmit power $P_{\mathrm{t}}$.
Fig. 6 illustrates the sum-rate performance versus transmit power for systems employing reconfigurable and conventional antennas. Specifically, the proposed tri-hybrid beamforming architecture with reconfigurable antennas and a dynamic-subarray scheme is denoted as “RA, DS”. For comparison, hybrid beamforming using conventional antennas with a dynamic-subarray is labeled “CA, DS”, and the fixed-subarray architectures with reconfigurable and conventional antennas are denoted as “RA, FS” and “CA, FS”, respectively. Performance is evaluated for systems with $N_{\mathrm{t}}=32$ and $N_{\mathrm{t}}=16$ antennas. As shown in Fig. 6, the proposed “RA, DS” algorithm consistently outperforms its competitors. In particular, reconfigurable antennas yield a power gain of approximately 7 dB compared to conventional antennas, while the dynamic-subarray scheme provides an additional gain of about 2 dB over the fixed-subarray approach. These results confirm that the tri-hybrid beamforming architecture achieves superior communication performance by virtue of its enhanced flexibility in radiation pattern selection in the EM domain and subarray selection in the RF domain.
@@
For the proposed tri-timescale tri-hybrid beamforming design, increasing the channel sampling rate in the medium-timescale can enhance sum-rate performance by enabling more frequent updates of channel statistics for multi-path rays and scattering cluster cores. However, this improvement comes at the cost of higher pilot overhead. To quantitatively evaluate the impact of the channel sampling rate on both system performance and pilot overhead, we gradually increase the number of frames per super-frame in the medium-timescale, effectively increasing the channel sampling rate, as illustrated in Fig. 7. By analyzing the pilot overhead expression derived in Section III-E and observing the trend in Fig. 7, we conclude that the proposed tri-timescale scheme significantly reduces pilot overhead compared to the real-time scheme, which requires high-dimensional CSI estimation in every time-slot. More importantly, as the number of frames increases, the sum-rate performance of the tri-timescale beamforming scheme steadily improves and eventually saturates at a satisfactory level, albeit with a monotonic increase in channel estimation overhead. In particular, setting the number of frames to $T_{\mathrm{M}}=50$ achieves a favorable balance between performance and overhead, a configuration adopted in the simulations.
Figure 8: The impact of channel estimation configurations on sum-rate performance.
The channel estimation algorithms in both the EM and RF domains are critical components of the proposed tri-timescale framework. Fig. 8 illustrates the impact of different channel estimation configurations on sum-rate performance under various system settings. The results indicate that, regardless of the number of propagation paths $L_{\mathrm{c}}$, the sum-rate achieved with estimated CSI is highly comparable to that with perfect CSI. This finding underscores the accuracy and robustness of our channel estimation algorithms within the proposed framework, even under complex propagation conditions. Moreover, even with a reduced angular resolution $N_{\mathrm{Beam}}$, the channel characteristics are reliably captured, and acceptable sum-rate performance is achieved, further validating the effectiveness and reliability of the proposed approach.
Figure 9: The sum-rate versus the distance and angle of UE.
@@
Finally, we validate the effectiveness of the proposed scheme by evaluating the impact of user location on the overall performance of the tri-hybrid beamforming with reconfigurable antennas (RA) and the hybrid beamforming with conventional antennas (CA). Specifically, Fig. 9 (a) depicts the sum-rate performance as a function of the nominal distance, which represents the distance between the BS and the UEs, while Fig. 9 (b) illustrates the sum-rate variation with respect to the nominal angle, which represents the user’s angular position relative to the antenna array. From Fig. 9 (a), we observe that as the nominal distance increases, the performance gain offered by pattern-reconfigurable antennas gradually diminishes. This is primarily due to the dominant effect of path loss, which attenuates signal power, reducing the relative impact of radiation pattern variations across different angles. In contrast, Fig. 9 (b) reveals that the RA-based scheme maintains robust performance across a wide range of user angles by dynamically adjusting the pattern orientations as the user’s position deviates from the optimal direction. In comparison, the CA-based scheme achieves satisfactory performance only when the user is positioned perpendicular to the array, where the antenna gain is maximized. These results underscore the robustness and versatility of the proposed tri-hybrid beamforming approach, demonstrating that dynamic pattern reconfiguration not only compensates for variations in user location but also substantially enhances overall system performance and reliability across a wide range of deployment scenarios.
@@ V Conclusions
This paper presents a novel tri-hybrid beamforming architecture that integrates pattern-reconfigurable antennas with analog and digital beamforming. By leveraging an innovative tri-timescale beamforming framework, the proposed approach optimizes the radiation beamformer over a long-timescale, the analog beamformer over a medium-timescale, and the digital beamformer over a short-timescale. This hierarchical design effectively balances system performance with practical constraints, such as channel estimation overhead and computational complexity. Extensive simulation results validate the superiority of the proposed scheme over conventional hybrid beamforming architectures, highlighting its potential in future wireless communication systems.

### v2

@@ III-E.2 Complexity Analysis
Next, we provide a brief analysis of the algorithmic complexity. The primary computational complexity lies in calculating the EM domain pattern selection matrix $\mathbf{S}_{\mathrm{EM}}$ in (36) over the long-timescale, and determining the phase-shift selection matrix $\mathbf{S}_{\mathrm{RF}}$ in (56) over the medium-timescale. Specifically, the computational complexity of solving problem (36) in each super-frame over the long-timescale with an $N_{\mathrm{t}}P$-dimensional variable, is of order $\mathcal{O}\{N_{\mathrm{t}}^{3.5}P^{3.5}\}$ [38]. Similarly, the computational complexity of solving problem (56) in each frame during the medium-timescale with an $N_{\mathrm{t}}N_{\mathrm{RF}}2^{B}$-dimensional variable is of order $\mathcal{O}\{N_{\mathrm{t}}^{3.5}N_{\mathrm{RF}}^{3.5}2^{3.5B}\}$.
@@ III-E.3 Pilot Overhead Analysis
The pilot overhead in the proposed scheme primarily stems from channel sampling at the medium-timescale and short-timescale stages. In the medium-timescale, CS-based channel estimation requires $I$ pilot symbols $(N_{\mathrm{RF}}I\gg L)$ for each UE to obtain the RF domain high-dimensional statistical channel sample $\mathbf{h}_{\mathrm{RF},k}^{(t_{\mathrm{M}})}$, as described in Section III-C. In the short-timescale, $N_{\mathrm{RF}}K$ pilot symbols are needed for estimating the BB domain low-dimensional effective channel $\mathbf{H}_{\mathrm{e}}^{(t_{\mathrm{S}})}$, as described in Section III-D. Therefore, the total pilot overhead across the channel statistics coherence time is expressed as $(KI+KT_{\mathrm{S}})T_{\mathrm{M}}T_{\mathrm{L}}$. This is notably less than that of the real-timescale beamforming approaches, which requires $(KI+K)T_{\mathrm{S}}T_{\mathrm{M}}T_{\mathrm{L}}$.
III-E.4 Initialization
During the initialization phase, we set all antenna patterns to point toward the boresight direction for the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$, adopt a random phase initialization for the analog beamformer $\mathbf{F}_{\mathrm{RF}}$, and employ a normalized identity matrix for the digital beamformer $\mathbf{F}_{\mathrm{BB}}$. These initial settings serve as a neutral starting point for the subsequent optimization.
@@ IV Simulation Results
In this section, we provide simulation results to demonstrate the effectiveness of the proposed tri-timescale tri-hybrid beamforming design algorithm with reconfigurable antennas in enhancing communication performance while ensuring hardware efficiency. Unless otherwise specified, the BS is equipped with $N_{\mathrm{t}}=32$ antennas and the center carrier frequency is set to $f_{\mathrm{c}}=28$ GHz. The low-resolution PS used in the RF domain has quantized phase controlled by $B=3$ bits. For simplicity and to focus on the tri-timescale design, we set $N_{\mathrm{RF}}=K$ in simulations and the framework applies to $N_{\mathrm{RF}}\geq K$ as well. In the EM domain, each pixel-based reconfigurable antenna enables the generation of $P=7$ distinct directional radiation patterns by adjusting the electronic switches between pixels [12], with $M=180$ uniform sampling points in the azimuth dimension [^3] The prototype reconfigurable antenna in [12] currently has an aperture exceeding the conventional half-wavelength spacing, introducing practical challenges such as mutual coupling and grating lobes, which we leave for future investigation. In this study, we adopt half-wavelength spacing to facilitate fair comparisons with conventional array designs. With ongoing advancements in reconfigurable antenna technology, recent studies and prototypes have convincingly demonstrated the feasibility of smaller apertures compatible with half-wavelength spacing, thereby addressing these practical concerns.. In this scenario, there are $K=2$ UEs, all operating under a scattering cluster channel with $C=2$ clusters, each of which consists of $L_{c}=3$ multi-path rays. The nominal AoDs $\overline{\theta}_{k,c}$ of these scattering clusters are randomly distributed within $\pi/2,\pi/2]$, while the angle spread is set to $\varsigma_{k,c}=\pi/36$. The large-scale path loss is expressed as $\alpha_{k,c}=10^{-\frac{C_{0}}{10}}(\frac{r_{k,c}}{D_{0}})^{-\kappa_{k,c}}$ with $\kappa_{k,c}\in[2.5,3]$ being the attenuation factor, $C_{0}=30$ dB, $D_{0}=1$m. The propagation distance $r_{k,c}$ of each scattering cluster is randomly distributed in $r_{k,c}\in[50\mathrm{m},100\mathrm{m}]$. The noise power of UEs is set as $\sigma_{k}^{2}=-70$dBm. The tri-timescale framework is structured with $T_{\mathrm{L}}=10$ super-frames for the long-timescale, $T_{\mathrm{M}}=50$ frames for the medium-timescale, and $T_{\mathrm{S}}=200$ time-slots for the short-timescale.
@@
Figure 5: Convergence performance of the radiation beamformer design and analog beamformer design.
Figure 6: The sum-rate versus the transmit power $P_{\mathrm{t}}$.
@@
Figure 7: The sum-rate versus the distance and angle of UE.
@@
Fig. 6 illustrates the sum-rate performance versus transmit power for systems employing reconfigurable antennas (RA) and conventional antennas (CA). The proposed tri-hybrid beamforming architecture with reconfigurable antennas and dynamic-subarray selection is labeled as “RA, DS”. For benchmarking purposes, fully-digital beamforming with reconfigurable antennas denoted as “RA, FD” serves as the upper performance bound and corresponds to a real-timescale optimization, while hybrid beamforming with fixed-subarray selection denoted as“RA, FS” represents the lower bound. Specifically, the “RA, DS and “RA, FS” configurations implement the proposed tri-timescale optimization framework, where EM domain, RF domain, and BB domain beamformers are optimized across distinct timescales. For comparison, we also include three schemes employing conventional antennas: “CA, DS”, “CA, FD”, and “CA, FS”. Among them, “CA, FD” corresponds to a real-timescale design, whereas “CA, DS” and “CA, FS” employ a conventional two-timescale optimization approach, optimizing analog beamformers over a long timescale and digital beamformers over a short timescale [28]-[30]. As observed in Fig. 6, the proposed “RA, DS” approach consistently outperforms other schemes. In particular, the use of reconfigurable antennas achieves approximately a 7 dB gain compared to conventional antennas. Furthermore, employing dynamic-subarray selection yields an additional gain of roughly 2 dB over fixed-subarray schemes. These results validate that the proposed tri-hybrid beamforming architecture, benefiting from adaptive EM domain radiation patterns and flexible RF domain subarray selections, significantly enhances overall communication performance.
(a) Long-timescale.
(b) Medium-timescale.
Figure 9: The performance evaluation of channel estimation in the long-timescale and medium-timescale.
Figure 10: The sum-rate versus the number of angular sampling points $M$.
@@
For the proposed tri-timescale tri-hybrid beamforming design, increasing the channel sampling rate in the medium-timescale can enhance sum-rate performance by enabling more frequent updates of channel statistics for multi-path rays and scattering cluster cores. However, this improvement comes at the cost of higher pilot overhead. To quantitatively evaluate the impact of the channel sampling rate on both system performance and pilot overhead, we gradually increase the number of frames per super-frame in the medium-timescale, effectively increasing the channel sampling rate, as illustrated in Fig. 8. By analyzing the pilot overhead expression derived in Section III-E and observing the trend in Fig. 8, we conclude that the proposed tri-timescale scheme significantly reduces pilot overhead compared to the real-time scheme, which requires high-dimensional CSI estimation in every time-slot. More importantly, as the number of frames increases, the sum-rate performance of the tri-timescale beamforming scheme steadily improves and eventually saturates at a satisfactory level, albeit with a monotonic increase in channel estimation overhead. In particular, selecting the number of frames as $T_{\mathrm{M}}=50$ achieves an optimal balance between performance and overhead, and this configuration is therefore employed in our simulations.
Figure 11: The sum-rate versus the number of antennas and UEs.
@@
Next, we validate the effectiveness of the proposed scheme by evaluating the impact of user location on the overall performance of the tri-hybrid beamforming with reconfigurable antennas and the hybrid beamforming with conventional antennas. Specifically, Fig. 7 (a) depicts the sum-rate performance as a function of the nominal distance, which represents the distance between the BS and the UEs, while Fig. 7 (b) illustrates the sum-rate variation with respect to the nominal angle, which represents the user’s angular position relative to the antenna array. From Fig. 7 (a), we observe that as the nominal distance increases, the performance gain offered by pattern-reconfigurable antennas gradually diminishes. This is primarily due to the dominant effect of path loss, which attenuates signal power, reducing the relative impact of radiation pattern variations across different angles. In contrast, Fig. 7 (b) reveals that the RA-based scheme maintains robust performance across a wide range of user angles by dynamically adjusting the pattern orientations as the user’s position deviates from the optimal direction. In comparison, the CA-based scheme achieves satisfactory performance only when the user is positioned perpendicular to the array, where the antenna gain is maximized. These results underscore the robustness and versatility of the proposed tri-hybrid beamforming approach, demonstrating that dynamic pattern reconfiguration not only compensates for variations in user location but also substantially enhances overall system performance and reliability across a wide range of deployment scenarios.
The channel estimation algorithms in both the EM and RF domains are critical components of the proposed tri-timescale framework. As shown in Fig. 9 (a), the normalized APS obtained by the Capon method is depicted as a function of the angle. Two dominant peaks can be clearly observed, corresponding to the scattering clusters in the angular domain. The estimated nominal angles, marked by red dots, are well aligned with the peaks of the Capon APS, indicating accurate cluster angle estimation in the long timescale. The normalized mean square error (NMSE) performance of channel estimation is evaluated versus SNR for different schemes. The CS-based channel estimation using dictionary with pattern information consistently outperforms its counterpart without pattern information, demonstrating the benefit of exploiting pattern-aware dictionaries in the medium-timescale channel estimation process.
In Fig. 10, we investigate how the number of angular sampling points $M$ affects the achievable sum-rate. It is evident that the system sum-rate monotonically increases with $M$, since a finer angular grid provides a more accurate discretization of continuous radiation patterns, thus improving the matching precision between radiation gains and user angles. Nevertheless, the incremental benefit gradually becomes negligible as $M$ increases beyond a certain point (around $M=60$ for perfect CSI and $M=180$ for estimated CSI), indicating limited value in excessively dense angular sampling. Specifically, after comprehensive consideration, the choice of $M=180$, adopted in earlier simulations, achieves a favorable balance between performance improvement and computational complexity, effectively capturing most performance gains without incurring excessive complexity.
To further examine the scalability of the proposed tri-hybrid beamforming architecture, Fig. 11 illustrates the achievable sum-rate performance under varying numbers of BS antennas and UEs. The results in Fig. 11 (a) clearly underscore the hardware efficiency and substantial array-size reductions achievable with reconfigurable antennas, highlighting their considerable potential for supporting energy-efficient next-generation wireless systems without compromising communication performance. Additionally, as depicted in Fig. 11 (b), the proposed tri-hybrid beamforming consistently outperforms both baseline hybrid beamforming and conventional antenna schemes across all considered user loads, demonstrating robust scalability. These comprehensive simulation results thus strongly validate the practicality and scalability of the proposed tri-hybrid architecture for future large-scale multi-user wireless deployments.
@@ V Conclusions
This paper presents a novel tri-hybrid beamforming architecture that integrates pattern-reconfigurable antennas with analog and digital beamforming. By leveraging an innovative tri-timescale beamforming framework, the proposed approach optimizes the radiation beamformer over a long-timescale, the analog beamformer over a medium-timescale, and the digital beamformer over a short-timescale. This hierarchical design effectively balances system performance with practical constraints, such as channel estimation overhead and computational complexity. Numerical results confirmed that the proposed design achieves around 7 dB sum-rate improvement, while reducing pilot overhead by nearly an order of magnitude compared with real-time joint beamforming schemes. The demonstrated scalability and performance gains underscore the practical feasibility and effectiveness of our approach for future large-scale wireless systems. Future work will address practical hardware aspects and extend this framework to wideband and near-field scenarios.
Future research directions include addressing practical hardware challenges associated with reconfigurable antennas, such as severe mutual coupling effects resulting from half-wavelength spacing. Incorporating multi-port circuit theory to model and mitigate mutual coupling through joint radiation pattern and hybrid beamforming design is a promising pathway. Furthermore, larger antenna spacing can introduce strong grating lobes in practice. To address this issue, it will be essential to explore sparse or non-uniform array configurations, irregular sampling methods, and integrated radiation pattern optimization schemes. Investigating these aspects will further extend the practical applicability and deepen the theoretical understanding of tri-hybrid beamforming architectures for realistic deployment scenarios.